an open box is made from a 40-cm by 80-cm piece of tin by cutting a square from each corner and folding up the edges. The area of the resulting base is 2964 cm^2. What is the length of the sides of the squares
Answer by ikleyn(52754)

.
an open box is made from a 40-cm by 80-cm piece of tin by cutting a square from each corner
and folding up the edges. The area of the resulting base is 2964 cm^2.
What is the length of the sides of the squares
~~~~~~~~~~~~~~~~~~~~~~

After cutting the squares from the corners and folding,
the dimensions of the base are (40-2x) cm and (80-2x) cm,
where x is the side of each square.


Now you write this quadratic equation for the base area

    (40-2x)*(80-2x) = 2964  square centimeters.


Then you simplify it step by step

    (20-x)*(40-x) = 741

     800 - 40x - 20x + x^2 = 741

     x^2 - 60x + 59 = 0

     (x-59)*(x-1) = 0

and find the roots  59 and 1.


The root 59 is too big to be the side of the cut square,
so the only solution to the problem is 1 cm for the side of the squares.


CHECK.  (40-2*1)*(80-2*1) = 2964 cm^2, the area of the base.  ! correct !

Solved.